### v2

@@ 4 Proposed Method
This section provides a detailed explanation of the proposed Retrieval-enhanced Prompt Tuning for VLPT. RePrompt enhances prompt tuning by utilizing the retrieval of relevant information from the training dataset. Given an input image $x$, RePrompt retrieves $k$ potentially useful entries from the retrieval database. Each entry is associated with a corresponding representation $z$ and a label (the construction of the retrieval database is described in Section 4.1). The retriever uses embedding similarity to find relevant downstream knowledge in the retrieval database. We then condition on both the retrieved set and the original input $x$ to generate the output through retrieval-enhanced visual prompt and retrieval-based adapter. Figure 2 illustrates the workflow of RePrompt. Specifically, Section 4.1 describes the process of retrieving entries most relevant to the input query. Subsequently, Section 4.2 details the prompt learner that integrates the query and retrieved knowledge into the prompt learning process. In Section 4.3, we discuss the final prediction in the inference process using the $k$NN-based probability from the retrieval-based adapter. Additionally, this section describes how non-parametric KNN is used between the query and the database to obtain probability distributions, while dynamically controlling the strength of the retrieval mechanism.
4.1 Retrieval module
Our Retrieval module consists of two steps: (1) building the database, and (2) retrieval.
@@ 4.1 Retrieval module
Retrieval database. The retrieval database is constructed by extracting features from the few-shot training dataset $\mathcal{D}$. Specifically, the database contains $\left|\mathcal{D}\right|$ key-value pairs $\left(\boldsymbol{k}_{i},\boldsymbol{v}_{i}\right)$. Each key $\boldsymbol{k}_{i}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ represents the frozen training image representation extracted using $\boldsymbol{e}_{\mathrm{R}}$. Each key-value pair comprises two components: the label $\boldsymbol{y}\in\mathbb{N}^{+}$ and the image representation $\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$. The retrieval database acts as a repository of robust tokens that are adaptable to changes in downstream tasks.
@@
Previous studies retrieval_long_tail ; retrieval_pretraining ; RETRO_tokens have demonstrated the effective performance improvements achieved by expanding the retrieval database. Direct use of external data can disrupt the few-shot task setting. To address this issue, we employ stable diffusion techniques rombach2022high to generate additional training data for various categories. This approach allows us to expand the training data without requiring additional human effort for data collection and annotation. A straightforward template “a photo of a [CLASS]” is utilized as the text prompt. Subsequently, we employ CLIP to filter the top-$\hat{K}$ images of the highest quality for each category. The selected images are added to the original retrieval database, expanding it with new samples and resulting in a total of ($K+\hat{K}$) training images across $C$ categories. To ensure a fair comparison, synthesized data are not included in the training datasets used for the main experiments. In Table 5, we present the performance improvement achieved by incorporating synthesized data, demonstrating an average precision of 84.19% for the 16-shot scenario.
Figure 4: Examples of the retrieval mechanism using various encoders $e_{\mathrm{R}}$ on ImageNet under the 16-shot setting. ”ViT-B/16 (Image encoder of CLIP)” indicates that $e_{\mathrm{R}}$ utilizes the ViT-B/16 model as trained by CLIP. ”ViT-B/16 (Timm)” denotes that $e_{\mathrm{R}}$ employs the ViT-B/16 model as trained by Timm with full supervision. ”Top $k$” refers to the $k$ nearest neighbors, and ”IPS” represents the inner product similarity. Employing the image encoder of CLIP as $e_{\mathrm{R}}$ enhances the retrieval of visually similar and more accurate images, particularly for challenging samples such as ”velvet fabric”.
@@
Effective and efficient retrieval. As shown in Figure 2, the retrieval database is represented as a matrix $\boldsymbol{D}\in\mathbb{R}^{\left|\mathcal{D}\right|\times d}$, which serves as a fast approximate $k$-NN of examples. When presented with a query image $\boldsymbol{x}_{\mathrm{q}}$, we utilize the function $e_{\mathrm{R}}$ to map it to a corresponding vector $\boldsymbol{z}_{\mathrm{q}}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{\mathrm{q}})$. Using the query vector $\boldsymbol{z}_{\mathrm{q}}$, we retrieve its approximate $k$-nearest neighbors with representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k}$ from the matrix $\boldsymbol{D}$ based on cosine similarity.
To demonstrate the impact of our retrieval mechanism and the influence of different $e_{\mathrm{R}}$ functions, we include visualizations of retrieved images corresponding to query samples in Figure 4. The visualization results demonstrate the retrieval’s capability to locate similar samples, with the image encoder of CLIP proving to be an effective $e_{\mathrm{R}}$ in identifying more accurately matched images. To ensure an efficient retrieval process, we employ FAISS faiss for querying the database.
4.2 Retrieval-enhanced visual prompt
@@
The retrieval module receives a query vector $\boldsymbol{z}_{\mathrm{q}}$ derived from a raw image $\boldsymbol{x}_{\mathrm{q}}$ and performs a lookup operation in the matrix $\boldsymbol{D}$ to obtain the top $k_{\mathrm{re}}$ most similar candidates. The corresponding representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ retrieved from the database are integrated into the image encoder to enhance the visual prompts. Additional fusion vectors $\boldsymbol{z}_{\mathrm{f}}\in\mathbb{R}^{d}$ RETROPROMPT are generated by intuitively aggregating $k_{\mathrm{re}}$ neighboring representations based on their similarity, as follows:
@@
As depicted at the top of Figure 3, we randomly initialize $J$ visual prompt sets, $\boldsymbol{P}^{1}_{\mathrm{I}},\dots,\boldsymbol{P}^{J}_{\mathrm{I}}$. Subsequently, $J$ retrieval-enhanced convolution (REConv) blocks process $J$ inputs, $\hat{\boldsymbol{i}}_{1},\dots,\hat{\boldsymbol{i}}_{J}$ (by replicating $\hat{\boldsymbol{i}}$ $J$ times), to generate dynamic prompts. These prompts are combined separately with the $J$ visual prompt sets and inserted into the input sequences of the first $J$ layers of the visual branch:
@@
The remaining $12-J$ layers process learnable prompts without enhancement, and learning these prompts reverts to conventional visual prompt tuning. Furthermore, we discuss the choice of insertion depth in subsection 6.2.
To effectively and efficiently fuse the retrieved representations, we propose a retrieval-enhanced convolution block (REConv). As illustrated at the bottom of Figure 3, a REConv block comprises three convolution layers: two $1\times 1$ convolutions, which individually reduce and scale the channel dimensionality, and a $3\times 3$ convolution positioned between the two $1\times 1$ convolutions. Before these convolution layers, we reshape the 1D token sequence structure of visual prompts into a 2D matrix structure. REConv blocks process the input $\hat{\boldsymbol{i}}$ in parallel to generate dynamic prompts, which can be formulated as
@@
4.3 Retrieval-based adapter
@@
Relation to Tip-Adapter. Tip-Adapter constructs an adapter using a key-value cache model from the few-shot training set. The cache model stores prior knowledge encoded in CLIP and facilitates image classification through retrieval in the image embedding space. Tip-Adapter is considered as an auxiliary classifier that leverages image-image similarity. While both approaches aim to improve performance using retrieval, our paper highlights two key distinctions:
  1. (i)
    The features extracted by the CLIP image encoder with retrieval-enhanced prompts are more adaptive to downstream tasks compared to the features obtained by the original CLIP image encoder in Tip-Adapter. As a result, prompt-tuned query features can be utilized to compute improved image-image and image-text similarity matrices for classification.
  2. (ii)
    We extract potential visual context from the cache model and integrate it into the model in the form of prompts, rather than considering the cache model solely as an adapter for classification assistance.
4.4 Retrieval-guiding training
The $k$-nearest neighbor ($k$NN) algorithm primarily focuses on approximating the neighborhoods of query instances knn_classifier . It is intuitive to leverage $k$NN classification results as prior knowledge to guide RePrompt’s focus on hard examples during the training process. Hard samples usually refer to atypical samples with low confidence. To compute a local probability distribution, we limit the number of samples within the retrieved neighbor set $\boldsymbol{K}\subseteq\boldsymbol{D}$, where $k_{rc}\neq\left|\boldsymbol{K}\right|$, as follows:
@@ 4.4 Retrieval-guiding training
The probability $\boldsymbol{p}_{k\mathrm{NN}}$ represents the confidence of classifying the query instance $\boldsymbol{x}_{\mathrm{q}}$ into specific categories. The negative log-likelihood value of $\boldsymbol{p}_{k\mathrm{NN}}$, similar to Focal Loss focal_loss , serves as the adjustment factor $\boldsymbol{p}_{\mathrm{t}}=-\log\left(\boldsymbol{p}_{k\mathrm{NN}}\right)$. This adjustment factor modifies the relative loss of pseudo-correct or pseudo-error samples distinguished by $k$NN, thereby reweighting the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$. The final loss is formulated as follows:
@@
where $\gamma$ is a scaling factor. We set $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n},\boldsymbol{n}\in\mathbb{N}^{+}$ RETROPROMPT , which adopts a similar loss for augmenting the model performance on the few-shot learning task of NLP. In few-shot experiments, $\boldsymbol{n}$ can take values of $1,2,4,8,16$ to accommodate different few-shot settings. We discuss the choice of $n$ in Table 10.
5 Experiments
5.1 Few-shot classification
Datasets. The RePrompt model is evaluated on 11 image classification datasets, covering various tasks such as object recognition (ImageNet1k imagenet_paper and Caltech101  Caltech101_paper ), fine-grained object recognition (Oxford Pets oxford_pest_paper , Flowers102 oxford_flowers_paper , FGVC Aircraft fgvc_aircraft_paper , Stanford Cars Stanford_cars_paper , and Food101 food101_paper ), sense recognition (SUN397 sun397_paper ), remote sensing recognition (EuroSAT eurosat_paper ), texture recognition (DTD DTD_paper ), and action recognition (UCF101 ucf101_paper ). we construct 1,2,4 ,8,16-shot training sets with full categories and the entire test set.
Baseline. We compare our proposed RePrompt with these existing prompt-based approaches: CoOP CoOP , VPT-Deep VPT , Maple MaPLe , PromptSRC promptsrc , Tip-Adapter tip-adapter , Tip-Adapter-F tip-adapter , and VLPT. (1) CoOP learns the context prompt concatenated with $\left[\mathrm{CLASS}\right]$ as the input of text encoder. (2) VPT-Deep incorporates learnable visual prompts into each transformer layer of the visual encoder. (3) VLPT jointly optimizes prompts across different modality encoders, following the methodologies of CoOP and VPT-Deep. (4) Tip-Adapter constructs a cache-model-based adapter using few-shot training data. It is a model which performs few-shot predictions by using those retrieved samples in the end, simply by making some prototypes. (5) Tip-Adapter-F is the variant where the adapter is fine-tuned. (6) Maple learns prompts for the text and image branches simultaneously, rather than a separate side. (7) PromptSRC uses the clip’s original features to regularize the prompt learning.
Training details. Prompt tuning, originally introduced in NLP, is extended to unified visual and language models by implementing it on a ViT model, specifically ViT-B/16, with 12 transformer layers similar to the text encoder. The hyper-parameters of the retrieval modules are set as follows: $k_{\mathrm{re}}$ is set to 7, and 9 random prompts are initialized. The first 7 layers of the vision transformer are equipped with retrieval-enhanced prompts. The prompt and adapter are fine-tuned using the AdamW optimizer with a learning rate of 1e-3 and an epsilon value of 1e-4. We adhere to the data preprocessing protocol of CoOP CoOP and freeze the parameters inherited from the pre-trained model during training.
@@ 5.1 Few-shot classification
Results. The performance of baseline approaches and our proposed RePrompt for few-shot image classification is presented in Figure 5. In general, parameter-efficient finetuning methods perform better than zero-shot classification (zero-shot CLIP) in scenarios with an average of over 11 datasets. RePrompt overall provides consistent improvements on most shots in comparison with all existing methods. RePrompt demonstrates significant performance improvement, particularly with $+5.87\%$ accuracy on DTD and $+9.04\%$ accuracy on Stanford Cars, based on VLPT. RePrompt substantially enhances performance on challenging datasets that contain a wide range of categories, such as ImageNet with 1000 classes and SUN397 with 397 classes. We also observe that Reprompt achieves fewer improvements on Oxford Pets and Food101. This could be attributed to the presence of noisy data and unique data distribution patterns in these datasets CoOP ; food101_paper ; UPT .
5.2 Cross-dataset evaluation
@@
We evaluate the cross-dataset generalization ability of RePrompt by learning on the 16-shot ImageNet setting and then transferring it directly to the remaining 10 datasets. Since this evaluation does not provide training images for other datasets, we remove the retrieval-based adapter module from this test. Table 3 shows the performance comparison between CoOP, CoCoOP, Maple, and PrompSRC. On the ImageNet source dataset, RePrompt achieves the best performance and demonstrates a strong generalization performance by surpassing recent SOTA methods in 5 datasets.
5.3 Domain generalization

### v3

@@ IV Proposed Method
This section provides a detailed explanation of the proposed Retrieval-enhanced Prompt Tuning for VLPT. RePrompt enhances prompt tuning by utilizing the retrieval of relevant information from the training dataset. Given an input image $x$, RePrompt retrieves $k$ potentially useful entries from the retrieval database. Each entry is associated with a corresponding representation $z$ and a label (the construction of the retrieval database is described in Section IV-A). The retriever uses embedding similarity to find relevant downstream knowledge in the retrieval database. We then condition both the retrieved set and the original input $x$ to generate the output through the retrieval-enhanced visual prompt and retrieval-based adapter. Fig. 2 illustrates the workflow of RePrompt. Specifically, Section IV-A describes the process of retrieving entries most relevant to the input query. Subsequently, Section IV-B details the prompt learner that integrates the query and retrieved knowledge into the prompt learning process. In Section IV-C, we discuss the final prediction in the inference process using the $k$NN-based probability from the retrieval-based adapter. Additionally, this section describes how non-parametric KNN is used between the query and the database to obtain probability distributions, while dynamically controlling the strength of the retrieval mechanism.
IV-A Retrieval Module
Our retrieval mechanism/process consists of two steps: (1) retrieval database construction, and (2) nearest neighbor retrieval.
@@ IV-A Retrieval Module
Retrieval database construction. The retrieval database is constructed by extracting features from the few-shot training dataset $\mathcal{D}$. Specifically, the database contains $\left|\mathcal{D}\right|$ key-value pairs $\left(\boldsymbol{k}_{i},\boldsymbol{v}_{i}\right)$. Each key $\boldsymbol{k}_{i}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ represents the training image representation encoded by the frozen image encoder $\boldsymbol{e}_{\mathrm{R}}$. In our method, values of each key-value pair comprise two components: the label $\boldsymbol{y}\in\mathbb{N}^{+}$ (one-hot encoding) and the image representation $\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ encoded by $\boldsymbol{e}_{\mathrm{R}}$. The retrieval database acts as a repository of robust tokens that are adaptable to changes in downstream tasks.
@@
Previous studies [16, 19, 17] have demonstrated the effective performance improvements achieved by expanding the retrieval database. Direct use of external data can disrupt the few-shot task setting. To ensure a fair comparison, synthesized data are not included in the retrieval database used for main experiments. In the ablation study, we employ a stable diffusion generation model [47] to generate additional training data to expand training data without requiring additional human effort for data collection and annotation. The synthesized images are added to the original retrieval database. In Table V, we present the performance improvement achieved by incorporating synthesized data, demonstrating an average precision of 84.19% for the 16-shot scenario.
@@
Nearest neighbor retrieval. As shown in Fig. 2, the retrieval database is represented as a matrix $\boldsymbol{D}\in\mathbb{R}^{\left|\mathcal{D}\right|\times d}$, which serves as a fast approximate $k$-NN of examples. Given a query image $\boldsymbol{x}_{\mathrm{q}}$, we utilize the encoder $e_{\mathrm{R}}$ to map it to a corresponding vector $\boldsymbol{z}_{\mathrm{q}}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{\mathrm{q}})$. Using the query vector $\boldsymbol{z}_{\mathrm{q}}$, we retrieve its approximate $k$-nearest neighbors with representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k}$ from the matrix $\boldsymbol{D}$ based on cosine similarity. The image encoder of CLIP is proved to be an effective $e_{\mathrm{R}}$ in identifying more accurately matched images. To ensure an efficient retrieval process, we employ FAISS [48] for querying the database.
IV-B Retrieval-enhanced Visual Prompting
@@
The retrieval module receives a query vector $\boldsymbol{z}_{\mathrm{q}}$ derived from a raw image $\boldsymbol{x}_{\mathrm{q}}$ and performs a lookup operation in the matrix $\boldsymbol{D}$ to obtain the top $k_{\mathrm{re}}$ most similar candidates. The corresponding representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ retrieved from the database are integrated into the image encoder to enhance the visual prompts. Additional fusion vectors $\boldsymbol{z}_{\mathrm{f}}\in\mathbb{R}^{d}$ [18] are generated by intuitively aggregating $k_{\mathrm{re}}$ neighboring representations based on their similarity, as follows:
@@
As depicted at the top of Fig. 3, we randomly initialize $J$ visual prompt sets, $\boldsymbol{P}^{1}_{\mathrm{I}},\dots,\boldsymbol{P}^{J}_{\mathrm{I}}$. Subsequently, $J$ retrieval-enhanced convolution (REConv) blocks process $J$ inputs, $\hat{\boldsymbol{i}}_{1},\dots,\hat{\boldsymbol{i}}_{J}$ (by replicating $\hat{\boldsymbol{i}}$ $J$ times), to generate dynamic prompts. These prompts are combined separately with the $J$ visual prompt sets and inserted into the input sequences of the first $J$ layers of the visual branch:
@@
The output token corresponding to the visual prompt is discarded. The remaining $12-J$ layers process learnable prompts without enhancement, and learning these prompts reverts to conventional visual prompt tuning.
We propose a retrieval-enhanced convolution block to effectively and efficiently fuse the retrieved representations (REConv). As illustrated at the bottom of Fig. 3, a REConv block comprises three convolution layers: two $1\times 1$ convolutions, which individually reduce and scale the channel dimensionality and a $3\times 3$ convolution positioned between the two $1\times 1$ convolutions. Before these convolution layers, we reshape the 1D token sequence structure of visual prompts into a 2D matrix structure. REConv blocks process the input $\hat{\boldsymbol{i}}$ in parallel to generate dynamic prompts, which can be formulated as
@@
IV-C Retrieval-based Adapter
@@
Relation to Tip-Adapter. Tip-Adapter constructs an adapter using a key-value cache model from the few-shot training set. The cache model stores prior knowledge encoded in CLIP and facilitates image classification through retrieval in the image embedding space. Tip-Adapter is considered as an auxiliary classifier that leverages image-image similarity. While both approaches aim to improve performance using retrieval, our paper highlights two key distinctions: (1) The features extracted by the CLIP image encoder with retrieval-enhanced prompts are more adaptive to downstream tasks compared to the features obtained by the original CLIP image encoder in Tip-Adapter. As a result, prompt-tuned query features can be utilized to compute improved image-image and image-text similarity matrices for classification. (2) We extract potential visual context from the cache model and integrate it into the model in the form of prompts, rather than considering the cache model solely as an adapter for classification assistance.
IV-D Retrieval-guided Training
The $k$-nearest neighbor ($k$NN) algorithm primarily focuses on approximating the neighborhoods of query instances[31]. It is intuitive to leverage $k$NN classification results as prior knowledge to guide RePrompt’s focus on hard examples during the training process. Hard samples usually refer to atypical samples with low confidence. To compute a local probability distribution, we limit the number of samples within the retrieved neighbor set $\boldsymbol{K}\subseteq\boldsymbol{D}$, where $k_{rc}\neq\left|\boldsymbol{K}\right|$, as follows:
@@ IV-D Retrieval-guided Training
The probability $\boldsymbol{p}_{k\mathrm{NN}}$ represents the confidence of classifying the query instance $\boldsymbol{x}_{\mathrm{q}}$ into specific categories. The negative log-likelihood value of $\boldsymbol{p}_{k\mathrm{NN}}$, similar to Focal Loss[49], serves as the adjustment factor $\boldsymbol{p}_{\mathrm{t}}=-\log\left(\boldsymbol{p}_{k\mathrm{NN}}\right)$. This adjustment factor modifies the relative loss of pseudo-correct or pseudo-error samples distinguished by $k$NN, thereby reweighting the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$. The final loss is formulated as follows:
@@
where $\gamma$ is a scaling factor. We set $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n},\boldsymbol{n}\in\mathbb{N}^{+}$[18], which adopts a similar loss for augmenting the model performance on the few-shot learning task of NLP. In few-shot experiments, $\boldsymbol{n}$ can take values of $1,2,4,8,16$ to accommodate few-shot settings. We discuss the choice of $n$ in Table X.
V Experiments
@@
V-A Few-shot Classification
Datasets. The RePrompt model is evaluated on 11 image classification datasets, covering various tasks such as object recognition (ImageNet1k [46] and Caltech101 [50]), fine-grained object recognition (Oxford Pets [51], Flowers102 [52], FGVC Aircraft [53], Stanford Cars [54], and Food101 [55]), sense recognition (SUN397 [56]), remote sensing recognition (EuroSAT [57]), texture recognition (DTD [58]), and action recognition (UCF101 [59]). we construct 1,2,4 ,8,16-shot training sets with full categories and the entire test set.
Baseline. We compare our proposed RePrompt with these existing prompt-based approaches: CoOP [6], VPT-Deep [11], Maple [13], PromptSRC [15], Tip-Adapter [21], Tip-Adapter-F [21], and VLPT. (1) CoOP learns the context prompt concatenated with $\left[\mathrm{CLASS}\right]$ as the input of text encoder. (2) VPT-Deep incorporates learnable visual prompts into each transformer layer of the visual encoder. (3) VLPT jointly optimizes prompts across different modality encoders, following the methodologies of CoOP and VPT-Deep. (4) Tip-Adapter constructs a cache-model-based adapter using few-shot training data. It is a model that performs few-shot predictions by using those retrieved samples in the end, simply by making some prototypes. (5) Tip-Adapter-F is the variant where the adapter is fine-tuned. (6) Maple learns prompts for the text and image branches simultaneously, rather than a separate side. (7) PromptSRC uses the clip’s original features to regularize the prompt learning.
Training details. Prompt tuning, originally introduced in NLP, is extended to unified visual and language models by implementing it on a ViT model, specifically ViT-B/16, with 12 transformer layers similar to the text encoder. The hyper-parameters of the retrieval modules are set as follows: $k_{\mathrm{re}}$ is set to 7, and 9 random prompts are initialized. The first 7 layers of the vision transformer are equipped with retrieval-enhanced prompts. The prompt and adapter are fine-tuned using the AdamW optimizer with a learning rate of 1e-3 and an epsilon value of 1e-4. We adhere to the data preprocessing protocol of CoOP and freeze the parameters inherited from the pre-trained model during training.
@@ V-A Few-shot Classification
Results. The performance of baseline approaches and our proposed RePrompt for few-shot image classification is presented in Fig. 4. In general, parameter-efficient finetuning methods perform better than zero-shot classification (zero-shot CLIP) in scenarios with an average of over 11 datasets. RePrompt overall provides consistent improvements on most shots in comparison with all existing methods. RePrompt demonstrates significant performance improvement, particularly with $+5.87\%$ accuracy on DTD and $+9.04\%$ accuracy on Stanford Cars, based on VLPT. RePrompt substantially enhances performance on challenging datasets that contain a wide range of categories, such as ImageNet with 1000 classes and SUN397 with 397 classes. We also observe that Reprompt achieves fewer improvements on Oxford Pets and Food101. This could be attributed to the noisy data and unique data distribution patterns in these datasets [6, 55, 12].
@@
Cross-dataset Evaluation. We evaluate the cross-dataset generalization ability of RePrompt by learning on the 16-shot ImageNet setting and then transferring it directly to the remaining 10 datasets. Since this evaluation does not provide training images for other datasets, we remove the retrieval-based adapter module from this test. Table I shows the performance comparison between CoOP, CoCoOP, Maple, and PrompSRC. On the ImageNet source dataset, RePrompt achieves the best performance and demonstrates a strong generalization performance by surpassing recent SOTA methods in 5 datasets. The lower performance on FGVC Aircraft, DTD, and UCF101 can be attributed to the different nature of their tasks, which require more specialized feature extraction. However, we adopt the same retrieval database sampled from ImageNet.
V-B Domain Generalization